OpenStudy (anonymous):

let \[p_{n}\] denote the n-th prime in ascending order. Use induction to prove the following proposition. \[p_{n} \le 2^{2^{n-1}}\] let \[p_{n}\] denote the n-th prime in ascending order. Use induction to prove the following proposition. \[p_{n} \le 2^{2^{n-1}}\] @Mathematics

myininaya (myininaya):

how do we know \[p_1 \le 2^{2^{1-1}}\] i'm not totally sure what p_1 means

OpenStudy (zarkon):

p_1=2

OpenStudy (zarkon):

the first prime number

myininaya (myininaya):

oh

myininaya (myininaya):

lol

OpenStudy (anonymous):

p_1 = 2 p_2 = 3 p_3 = 5 etc

myininaya (myininaya):

\[2 \le 2 \text{ is true!}\]

OpenStudy (anonymous):

right

myininaya (myininaya):

so let's assume it is true for some integer k now so we have \[p_k \le 2^{2^{k-1}} \]

myininaya (myininaya):

now we need to show it is true for n=k+1

OpenStudy (anonymous):

\[p_{k+1} \le 2^{2^{k}}\]

OpenStudy (anonymous):

I just can't figure out how to go from the inductive hypothesis to that

myininaya (myininaya):

\[2^{2^{(k+1)-1}}=2^{2^{k-1} \cdot 2^{1}}=2^{2^{k-1}} \cdot 2^2 \ge p_k \cdot 2^2=2^2p_k\]

myininaya (myininaya):

thinking ...

myininaya (myininaya):

so we just showed \[2^2 p_k \le 2^{2^{(k+1-1)}}\] \[p_k \le \frac{2^{2^{k+1-1}}}{2^2}\] and now i'm stuck

OpenStudy (zarkon):

\[2^2^k=2^{2^{k-1}\cdot 2}=2^{2^{k-1}\cdot2^{2^{k-1}\]

OpenStudy (zarkon):

not displaying nice for me

myininaya (myininaya):

\[2^{2^k}=2^{{2^{k-1}} \cdot 2}=2^{2^{k-1}}\cdot2^{2^{k-1}}\]

OpenStudy (zarkon):

yes

OpenStudy (zarkon):

\[2^{2^k}=2^{{2^{k-1}} \cdot 2}=2^{2^{k-1}}\cdot2^{2^{k-2}2}=2^{2^{k-1}}\cdot2^{2^{k-2}}\cdot2^{2^{k-2}}=\cdots\]

OpenStudy (zarkon):

\[=\prod_{n=0}^{k}2^{2^n}\]

OpenStudy (anonymous):

what is that?

OpenStudy (zarkon):

product

OpenStudy (anonymous):

oh so it's like the sigma notation

myininaya (myininaya):

\[\text{ example of that notation \in use }\prod_{i=1}^{4}i=1 \cdot 2 \cdot 3 \cdot 4\]

OpenStudy (zarkon):

k-1 not k

myininaya (myininaya):

yes except you multiply

OpenStudy (anonymous):

so how does that help in this case?

OpenStudy (zarkon):

\[=\prod_{n=0}^{k-1}2^{2^n}\]

myininaya (myininaya):

\[p_1 \cdot p_2 \cdot \cdot \cdot p_k \le =\prod_{n=0}^{k-1}2^{2^n}\]

OpenStudy (zarkon):

up to k-1

myininaya (myininaya):

i put two many equal signs i copy what you have and didn't change it

OpenStudy (zarkon):

no ...it is k

myininaya (myininaya):

too*

OpenStudy (anonymous):

I'm getting lost

OpenStudy (asnaseer):

I think @Zarkon means:\[p_1 \cdot p_2 \cdot \cdot \cdot p_{k-1} \le =\prod_{n=0}^{k-1}2^{2^n}\]

myininaya (myininaya):

on yeah he does

OpenStudy (zarkon):

\[p_1 \cdot p_2 \cdot \cdot \cdot p_k\geq p_{k+1}\]

OpenStudy (zarkon):

I was wrong...it is up to k

myininaya (myininaya):

lol

OpenStudy (asnaseer):

:-)

myininaya (myininaya):

zarkon is wrong?

OpenStudy (anonymous):

is that the product of those primes?

OpenStudy (zarkon):

\[p_1 \cdot p_2 \cdot \cdot \cdot p_{k} \le \prod_{n=0}^{k-1}2^{2^n}\]

OpenStudy (asnaseer):

the Universe will soon end --- how can Zarkon make a MISHTAKE??? ;-)

myininaya (myininaya):

its horrifying

myininaya (myininaya):

zarkon use to be alien looking i wonder if that something to with it

OpenStudy (anonymous):

I've gotten very confused lol

myininaya (myininaya):

\[p_1 \le 2^{2^{1-1}}\] right?

myininaya (myininaya):

\[p_2 \le 2^{2^{2-1}}\] right?

OpenStudy (zarkon):

\[p_{k+1}\le p_1 \cdot p_2 \cdot \cdot \cdot p_{k} \le \prod_{n=0}^{k-1}2^{2^n}=2^{2^k}\]

myininaya (myininaya):

\[p_3 \le 2^{2^{3-1}}\]

OpenStudy (anonymous):

yes since \[3\le32\] I think

myininaya (myininaya):

... \[p_{k-3} \le 2^{2^{(k-3)-1}}\] \[p_{k-2} \le 2^{2^{(k-2)-1}}\] \[p_{k-1} \le 2^{2^{(k-1)-1}}\] \[p_{k} \le 2^{2^{k-1}}\]

myininaya (myininaya):

so since all those inequalities are true then \[p_1 \cdot p_2 \cdot \cdot \cdot p_{k} \le \prod_{n=0}^{k-1}2^{2^n} \]

OpenStudy (anonymous):

ok hold on, we started with \[p_{n} \le 2^{2^{n-1}}\] proved it for when n=2 then assumed \[p_{k} \le 2^{2^{k-1}}\] was true and needef to get to \[p_{k+1} \le 2^{2^{k}}\] what happens after this?

myininaya (myininaya):

what comes after \[p_{k+1} \le 2^{2^{k}}\] we are done you can put something after it like for all integers n>=1 such and such is

myininaya (myininaya):

is true*

OpenStudy (anonymous):

how do we get to that though

OpenStudy (anonymous):

somehow we find that\[2^{2^{n-1}}\] is equal to that product somehow

OpenStudy (asnaseer):

I think you can prove it like this:\[2^{2^k}=2^{2^{k-1+1}}=2^{2^{k-1}.2}=(2^{2^{k-1}})^2=p_k^2\]therefore:\[p_{k+1}=p_k^2\]and we have assumed that:\[p_k<=2^{2^{k-1}}\]therefore:\[p_k^2<=(2^{2^{k-1}})^2\]\[p_k^2<=2^{2^k}\]therefore:\[p_{k+1}<=2^{2^k}\]

OpenStudy (anonymous):

how do yo know that it is equal to (p_k)^2 in the first line shouldn't it be less than?

OpenStudy (asnaseer):

yes, sorry, my mistake in the 1st line.

OpenStudy (asnaseer):

\[2^{2^k}=2^{2^{k-1+1}}=2^{2^{k-1}.2}=(2^{2^{k-1}})^2>=p_k^2\]

OpenStudy (anonymous):

then in the second line how do you know that they are equal?

OpenStudy (anonymous):

all we know is that they are both \[\le 2^{2^{n-1}}\] right?

OpenStudy (asnaseer):

\[p_{k+1}>=p_k^2\]

OpenStudy (asnaseer):

hmmm - needs more thought...

OpenStudy (anonymous):

I don't think you can necessarily say that it is larger in that case

OpenStudy (asnaseer):

yes - I agree - it needs more thought - maybe Zarkon was giving a clue as to how to solve this - I will need to think over what he was trying to say.

OpenStudy (anonymous):

I am thoroughly confused as to what I should be doing

myininaya (myininaya):

\[2^{2^k}=2^{{2^{k-1}} \cdot 2}=2^{2^{k-1}}\cdot2^{2^{k-2}2}=2^{2^{k-1}}\cdot2^{2^{k-2}}\cdot2^{2^{k-2}}=\cdots\]

myininaya (myininaya):

do you understand this?

OpenStudy (anonymous):

yes I understand that

myininaya (myininaya):

do you understand this?

myininaya (myininaya):

ok so do you understand this \[2^{2^k}=2^{2^{k-1}} \cdot 2^{2^{k-2}} \cdot 2^{2^{k-3}} \cdots \cdot 2^{2^0}\]

OpenStudy (anonymous):

why is that a 0?

myininaya (myininaya):

do you understand this?

myininaya (myininaya):

\[2^{2^2}=2^{2^{2-1} \cdot 2} =2^{2^{2-1}} \cdot 2^{2^{2-2}2}=2^{2^{2-1}} \cdot 2^{2^0} \cdot 2^2\] \[=2^{2} \cdot 2^{2} \cdot 2^{0}\] what do you think about this?

myininaya (myininaya):

do you understand this?

myininaya (myininaya):

no wait i don't like my second line (i mean its right) but look at the first line

OpenStudy (anonymous):

hmm I think I get that

myininaya (myininaya):

do you understand this?

myininaya (myininaya):

\[2^{2^2}=2^{2^{2-1} \cdot 2} =2^{2^{2-1}} \cdot 2^{2^{2-2}2}=2^{2^{2-1}} \cdot 2^{2^0} \cdot 2^2 \] \[=2^{2^1} \cdot 2^{2^0} \cdot 2^{2^1}=2^{2^1} \cdot 2^{2^1} \cdot 2^{2^0}\]

myininaya (myininaya):

do you understand this?

myininaya (myininaya):

i'm just trying to show you the end pattern

OpenStudy (anonymous):

eventually your k will go to zero so you need to stop right?

myininaya (myininaya):

do you understand this?

myininaya (myininaya):

yes

OpenStudy (anonymous):

ok

myininaya (myininaya):

do you understand this?

myininaya (myininaya):

so that should piece in all the puzzle pieces now

myininaya (myininaya):

do you understand this?

myininaya (myininaya):

\[p_{k+1}\le p_1 \cdot p_2 \cdot \cdot \cdot p_{k} \le \prod_{n=0}^{k-1}2^{2^n}=2^{2^k}\] thats how zarkon got this

OpenStudy (anonymous):

where did that list of products of primes come from?

myininaya (myininaya):

do you understand this?

myininaya (myininaya):

hey brandon ignore the extra posts it is reposting stuff weird

OpenStudy (anonymous):

I'm not seeing extra posts here